OpenStudy (anonymous):

14 mclauss 0What is the equation in point-slope form of the line passing through (3, 6) and (-2, 1)? a. y + 6 = 1(x + 3) b. y + 1 = -1(x + 2) c. y - 6 = -1(x + 3) d. y - 6 = 1(x - 3)

OpenStudy (anonymous):

Should you attempt to plug x, y values in?

OpenStudy (anonymous):

First find slope m = ...

OpenStudy (anonymous):

i doint kno how

OpenStudy (anonymous):

m = Δy/ Δx = ...

OpenStudy (anonymous):

which do u think it is

OpenStudy (anonymous):

Of course I know which one, but then how can you do the next, next then next ones on your own?

OpenStudy (anonymous):

ikik could u show me how u get the answer cause i just have no idea

OpenStudy (anonymous):

m = Δy/ Δx = ( 6 -1) / ( 3 - (-2) ) = ....